OpenStudy (anonymous):

It is known that a bag of apples weighing 5.23 kg also weighs 11.53 pounds. From this information, derive an equation converting kg to pounds. Use it to calculate how much would a bag of oranges weighing 4.87 kg would weigh in pounds.

OpenStudy (amistre64):

the key is to find a constant of variation , cue thunder

OpenStudy (amistre64):

5.23 = 11.53 k; where k = the constant of variation ... k = 5.23/11.53 such that kg = (5.23/11.53)* lbs